OpenStudy (anonymous):

Write the slope-intercept of a line parallel to y= -2x-9 passing through the point (-2,3)

OpenStudy (anonymous):

The options are 2x+y=1 y=-2x-2 y=-2x-1 y=-2x+1

OpenStudy (anonymous):

so parallel lines have the same slope

OpenStudy (anonymous):

plug in the coordinates in the equation given

OpenStudy (anonymous):

it is 2x+y=1 plug in the coordinates

OpenStudy (jdoe0001):

what's the slope of say " y= -2x-9 " ?

OpenStudy (anonymous):

the slope intercept form of the equation is y=mx+b

Nnesha (nnesha):

the slope intercept form equation y= mx + b parallel lines have same slope so slope which you have in first equation is also be in 2nd then just plug your x and y values  plug all values then solve for b let me know if you have any question